OpenStudy (anonymous):

factor of: 16(a+b)^2-25(x-y)^2

OpenStudy (unklerhaukus):

\[m^2-n^2=(m-n)(m+n)\]

OpenStudy (anonymous):

how?

OpenStudy (unklerhaukus):

\[(m-n)(m+n)=m^2-nm+nm-n^2=m^2-n^2\]

OpenStudy (anonymous):

thank you, how about x^4+2x^2y^2+81y^4?

OpenStudy (anonymous):

please help

OpenStudy (unklerhaukus):

let x^2= u y^2= w

OpenStudy (anonymous):

i don't ge it

OpenStudy (anonymous):

*get

OpenStudy (unklerhaukus):

start like this \[x^4+2x^2y^2+81y^4=u^2+2uw+(9w)^2\]